a real estate agent believes that the average closing cost of purchasing a new home is $6500 over the purchase price . She selects 40 new home sales at random and finds that the average closing costs are $6600. the standard deviation of the population is $120. Test her belief at a=0.05
please help

Answer by stanbon(75887)

a real estate agent believes that the average closing cost of purchasing a new home is $6500 over the purchase price . She selects 40 new home sales at random and finds that the average closing costs are $6600. the standard deviation of the population is $120. Test her belief at a=0.05
------
Ho: u = 6500
Ha: u is not equal to 6500
-----------------------------
t(6600) = (6600-6500)/[120/sqrt(40)] = 5.2705
---
p-value = 2*P(t > 5.2705 when df = 39) = 0.0000053169..
--------------------
Conclusion: Since the p-value is less than 5%, reject Ho
at the 5% significance level.
===================================
Cheers,
Stan H.
===================================